a container holds a gas at a pressure of 55 kpa And temperature of -100.0°c. What will the pressure be when the temperature is increased to 200.0°c
1
Expert's answer
2020-05-24T03:01:56-0400

PV = nRT

nR / V = const

P = nRT / V = T × const

55 = 173 × const, const = 55 / 173 = 0.318

P = 0.318 × 473 = 150 kPa

Answer: P = 150 kPa.